If a gas is at 1.2atm, 25 degrees Celsius, and 10 liters what is the new volume if the pressure is changed to 2.8atm and the temperature to 35 degrees Celsius?
1
Expert's answer
2020-02-27T09:01:41-0500

4.43 liters

The general gas equation :P1V1/T1 = P2V2/T2. Where p1=1.2atm;

V1= 10 liters

T1= 273 + 25=298 K

P2=2.8 atm

T2= 273+35=308 K

V2 - ?


V2=(p1*V1*T2)/(T1*p2) ;

V2= 4,43 liters